# Topolectrical circuits – recent experimental advances and developments

### v1

@@ III.2.2 Circuit demonstrations of higher-dimensional NHSE phenomena
Typically, the NHSE drags all bulk states toward a boundary [231]. However, in two or more dimensions, Zou et al. [164] demonstrated that the NHSE can selectively pump topological modes while leaving the bulk unaffected, in what is known as the hybrid higher-order skin-topological (ST) effect [259]. The idea is that, in a multi-dimensional lattice, it is possible for topological modes under mixed boundary conditions to possess a spectral point gap, but not the non-topological modes. Further opening up the remaining boundary/s will then result in the NHSE only for the topological modes. In Ref. [164], the 2D network consists of a unit cell with four nodes, with asymmetric intra-cell couplings implemented using INICs [Fig. 4d1]. With suitable parameter tuning, the 2D and 3D circuit implementations [Figs. 4d2 and d3] exhibit skin-topological (ST) and skin-skin-skin (SSS) effects, respectively. These hybrid skin-topological modes arise from the interplay between topological and skin modes, and cannot exist without either. Notably, these ST modes can exhibit localized skin behavior despite the absence of net non-reciprocity, since the reciprocity is spontaneously broken by the topological modes taking unequal amplitudes in the two sub-lattices.
The hybrid skin-topological effect [259] has also inspired a scheme for switching on or off the NHSE by toggling the topological character of the perpendicular sub-chains, as was first proposed in the cold atom context [312]. When the switch is activated by adjusting the loss parameter, non-reciprocal pumping prevails, driven by non-reciprocity introduced by the switch. Zhang et al. [305] experimentally demonstrated this topological switching using their 3D circuit [Fig. 4e1]. The voltage dynamics across four different unit cells revealed distinct eigenmode distributions, such as skin-topo-skin and skin-topo-topo modes [Fig. 4e2, e3]. The absence or presence of skin modes corresponds to trivial or non-trivial topological modes, which are tuned through relays (model no: G6K-2F-Y) that switch resistors on or off between specific nodes.
@@ III.2.3 Circuit demonstrations of interacting NHSE phenomena
Despite being classical systems, electrical circuits can also simulate few-body interaction dynamics through mappings between many-body one-dimensional models and single-body multi-dimensional models [313, 48, 314, 315]. Besides their practical value, such mappings can also lend new physical insights [316, 317, 318], such as the interpretation of two-body repulsion as 2-dimensional boundary effects, and 3-body non-Hermitian interactions as feedback mechanism on a 2-dimension non-orthogonal non-Hermitian lattice whose GBZ description is notoriously subtle [244]. In particular, effective boundaries can be defined in the Hilbert space even when the lattice has PBCs along each direction. Reverse implementations of such mappings have also enabled small interacting quantum systems to simulate the physics in far larger multi-dimensional lattices [319, 320].
@@ III.2.4 Implementing NHSE without non-reciprocal couplings
Even though paradigmatic NHSE models i.e. the Hatano-Nelson model [321] almost always contain non-reciprocal (asymmetric) couplings, realizing the NHSE itself does not necessary require asymmetric couplings. In the so-called reciprocal non-Hermitian skin effect, a reciprocal system is designed such that it contains subsystems harboring equal and opposite NHSE, such that NHSE accumulation occurs in specific subsectors despite the system being reciprocal on the whole. This was first demonstrated by Hofmann et al. [127], who measured eigenmode localization at the edges of their 2D circuit [Fig. 4h1]. A diagonal imaginary hopping within the square plaquette makes the Hamiltonian globally non-Hermitian but still reciprocal; with a basis transform, it can be shown that it contains “hidden” equal and opposite NHSE chains at different transverse momentum sectors or effective flux. This imaginary hopping—realized as a resistor in their model [Fig. 4h2]—results in different hopping probabilities in the presence of this effective flux, resulting in unequal amplitudes in opposite directions [Fig. 4h3].
In 2 dimensions, a reciprocal 2D lattice can also exhibit nontrivial system size-dependent NHSE in the presence of higher-rank chirality [322], characterized by a non-conserved charge current [323]. First demonstrated in Ref. [322], a higher-rank chiral mode (rank-2) is also achieved by Zhu et al. [308] through gain/loss couplings that create an unbalanced chirality in the unidirectional chiral mode flow [Fig. 4i1]. This imbalance causes one specific mode to amplify and hence outlast others, leading to eventual mode localization. The long-lived rank-2 chiral mode dominates the system’s long-time dynamics and is observable in momentum-resolved measurements [Figs. 4i2, i3]. Remarkably, momentum-dependent input excitations lead to mode accumulation at specific driven locations in the lattice [Fig. 4i4]. This enables a versatile implementation of NHSE along a desired direction and position in a fully passive and reciprocal lattice, distinguishing it from the reciprocal skin effect presented in Ref. [127]. (Even in the absence of gain/loss and non-reciprocity in the lattice bulk, non-Hermiticity can emerge through the interaction Hamiltonian in an electromagnetically engineered metamaterial that is inversely designed from a TE lattice [324].)
@@ III.2.5 Demonstrating unconventional NHSE phenomena with long-range couplings
The extreme versatility of electrical circuit connections, which can directly couple any two desired nodes independent of their physical distance, make them ideal for implementing intrinsically long-ranged models. Since the NHSE is already highly non-local in that it leads to dramatic state accumulation in a spatially-distant edge, it is particularly susceptible to the effects of long-ranged hoppings that can change the overall real-space topology of the lattice. For instance, a single non-reciprocal impurity hopping in an otherwise homogeneous Hatano-Nelson chain is known to lead to enigmatic scale-free localization [325, 326, 327].
Recently, Guo et al. [167] introduced the new notion of scale-tailored localization (STL), distinct from conventional NHSE or Anderson localization [328]. The STL is a spectacular demonstration of how singularities in the skin localization length subtly interplay with hopping non-locality, as was implemented via unidirectional couplings with INIC circuit elements, and a long-range coupling controlled by switches [Fig. 4j1]. For this circuit, some of the admittance eigenvalues encircle the unit circle, while others are isolated within it [Fig. 4j2]. The spatial distribution of the corresponding STL eigenstates exhibits varying localization lengths that depend on the long-range coupling distance [Fig. 4j3].
III.2.6 Circuit setups for non-Hermitian sensing
@@
The primary interest in the NHSE stems from its strong dependence on local perturbations, such as a coupling between the two ends of a chain, which can drastically alter the OBC vs. PBC properties. Furthermore, a non-Hermitian topological lattice exhibits exponential sensitivity to its size due to the exponential growth of NHSE states, as formulated theoretically by Budich and Bergholtz [329], and popularized as non-Hermitian topological sensors [330]. This phenomenon was experimentally demonstrated by Yuan et al. [96] [Fig. 5a1]. The shift in resonance peak frequency occurs through two mechanisms: (i) control of the boundary coupling capacitance through the displacement and rotation of the copper electrodes [Fig. 5a2], and (ii) variations in the number of nodes in the circuit. The impedance resonances exhibit exponential behavior in response to circuit size [Fig. 5a3].
@@
In the previous non-Hermitian sensing experiments, the underlying mechanism is directional amplification, which results in a strongly in a non-local response. Indeed, exponential sensitivity intuitively requires a non-reciprocal non-Hermitian bulk, achieved through asymmetric couplings [332], which necessitate active elements such as op-amps in INICs [333, 334]. However, Zhang et al. [100] presented a fully passive circuit demonstrating a highly distinct impedance response dependent on the presence of a boundary coupling, even though the circuit uses only resistors as the non-Hermitian components.
The non-local impedance response in this passive circuit is due to the presence of “hidden” coupled Hatano-Nelson chains with opposite amplifications, as seen through a basis rotation, [Fig. 5c1]. After the basis transform, the effective momentum space circuit Laplacian at the resonant frequency is expressed as
@@
While this setup may seem to be just another demonstration of the reciprocal NHSE [127] (see Fig. 4i in Section III.2), the effective inter-chain coupling is set to be relatively weak compared to the intra-chain terms, such that the NHSE from both chains do not simply cancel in the intended experimental setup. However, due to the exponential sensitivity of the NHSE, they still do cancel for sufficiently long circuits, as a variation of the critical NHSE [296, 235, 297, 298] (refer to the discussion related to Fig. 4b in Section III.2). In this model, the circuit exhibits curious size-dependent isolated modes only in the presence of parasitic resistances [Fig. 5c4], when the OBC circuit is sufficiently long. Their spatial profiles are boundary-localized in a unique manner, as shown in the two rightmost plots in Fig. 5c4. Despite being robust against perturbations, such isolated modes do not appear to be protected by any conventional topological invariant, appearing at certain ranges of system sizes.
@@ IV.1 Non-linear phenomena in topolectrical circuits
Introducing non-linear dynamics into a system reveals intriguing phenomena that are unattainable in linear systems, such as higher harmonic generation [140, 335, 336], soliton wave propagation [337, 338], non-linear bands [339], non-linear interface modes [340] as well as various non-Hermitian interplays [341, 342, 343, 344, 345, 207, 346]. Specifically, in the presence of non-linear interactions, coherent collective dynamics can emerge among the autonomous units hosting non-linear oscillators. Of particular interest is the incorporation of a topological mechanism, which facilitates self-organized, protected boundary oscillations [347, 341, 340]. A common approach to realizing non-linear topological phenomena involves arranging non-linear oscillatory elements in a topological configuration, such as by dimerizing an array of oscillators in deference to the well-known SSH configuration.
@@
Lo et al. [350] recently presented a non-linear circuit experiment demonstrating NHSE that emerges depending on the intensity of the driving amplitude. Interestingly, this behavior arises not from asymmetric couplings but from Kerr-type non-linear dynamics [355, 356]. The circuit comprises back-to-back onsite non-linear varactors [Fig. 6h1] assembled into a circuit array inspired by the physics of Bogoliubov modes [357] [Fig. 6h2]. The non-linearity breaks pseudo-Hermiticity, leading to a spectral point gap that in turns gives rise to NHSE localization. The sideband modes [Fig. 6h3], distinct from higher-order harmonics, exhibit localized profiles that depend on the auxiliary drive frequency. Due to the Kerr non-linearity, the NHSE appear only when the input signal amplitude exceeds a certain threshold [Fig. 6h4]. This work highlights how non-linearity can induce the NHSE in classical electrical circuits.
@@
The complex emergent dynamical behavior in non-linear systems can sometimes pose significant challenges for achieving global synchronization [358]. However, Zhang et al. [351] recently demonstrated a Hatano-Nelson (HN) circuit hosting on-site non-linear Stuart-Landau oscillators [Fig. 6i1]. Due to the asymmetric couplings in the HN circuit, Hermiticity is broken, resulting in non-orthogonal eigenmodes. This facilitates more effective communication between modes, ultimately leading to global synchronization. This behavior manifests as collective voltage oscillations observed in the time domain [Fig. 6i2] and in the steady-state voltage profile [Fig. 6i3]. This study highlights how the NHSE can enable global synchronization in non-linear experimental circuits (also refer to Ref. [359]).
@@
A cornerstone of many non-linear systems is their ability to support chaotic dynamics [360, 361, 358, 352]. A paradigmatic chaotic system is the Chua oscillator which, unlike Stuart-Landau oscillators, exhibit irregular and chaotic orbits in some parameter regimes [362]. Mathematically described by the Lorenz system of differential equations, the Chua oscillator possesses a well-known circuit realization known as the Chua circuit [363, 364, 365]. It can be constructed with four passive components and one active component known as the Chua diode [366][Fig. 6j1], which provides the requisite non-linear negative resistance. This non-linearity is evident in its non-linear current-voltage () characteristic [Fig. 6j2], and typical chaotic behavior include the famed double-scroll chaotic attractor [Fig. 6j3].
@@
Of particular interest is how such chaotic behavior interplay with lattice band structure dynamics, such as when multiple non-linear chaotic oscillators are coupled to form an array. In Ref. [352], Sahin et al. investigated the consequences of coupling Chua oscillators in a topological SSH-like lattice structure [Fig. 6j4]. Since band topology is only rigorously defined in linear systems, one might intuitively expect the non-linearity, and especially its chaotic dynamics, to erode the topological localization. However, it turns out that topological edge localization persists robustly well into the non-linear regime, with the amplitudes of the chaotic scroll being very different at the bulk and edge oscillators [Fig. 6j5] only in the black parameter space region [Fig. 6j6] – the region extrapolated from the topological non-trivial parameter region defined in the linear limit. This extrapolation can be performed by tuning the - characteristic parameter of the Chua diode [Fig. 6j7], and explicitly showcase how the topological edge oscillation suppression robustly carries over from the linear into the non-linear regime.
@@
By externally perturbing the voltage oscillations by injecting current at the left edge node, one further observes that the chaotic oscillations, which are typically highly sensitive to external perturbations, persist robustly only when in the topologically non-trivial regime: In the regime connected to the trivial phase, the injected current reduces the chaotic double-scroll portraits to non-chaotic limit-cycle oscillations [Fig. 6j8]. However, in the non-trivial phase, the same current injection only disturbs the edge node oscillations, leaving the bulk oscillations mostly intact [Fig. 6j9]. This robustness is captured as a phenomenon termed “topologically protected chaos”, where chaotic dynamics counterintuitively become more robust by the remnants of topological localization deep inside the non-linear regime [352].

### v2

@@ III.2.2 Circuit demonstrations of higher-dimensional NHSE phenomena
Typically, the NHSE drags all bulk states toward a boundary [232]. However, in two or more dimensions, Zou et al. [164] demonstrated that the NHSE can selectively pump topological modes while leaving the bulk unaffected, in what is known as the hybrid higher-order skin-topological (ST) effect [259]. The idea is that, in a multi-dimensional lattice, it is possible for topological modes under mixed boundary conditions to possess a spectral point gap, but not the non-topological modes. Further opening up the remaining boundary(ies) will then result in the NHSE only for the topological modes. In Ref. [164], the 2D network consists of a unit cell with four nodes, with asymmetric intra-cell couplings implemented using INICs [Fig. 4d1]. With suitable parameter tuning, the 2D and 3D circuit implementations [Figs. 4d2 and d3] exhibit skin-topological (ST) and skin-skin-skin (SSS) effects, respectively. These hybrid skin-topological modes arise from the interplay between topological and skin modes, and cannot exist without either. Notably, these ST modes can exhibit localized skin behavior despite the absence of net non-reciprocity, since the reciprocity is spontaneously broken by the topological modes taking unequal amplitudes in the two sub-lattices.
The hybrid skin-topological effect [259] has also inspired a scheme for switching on or off the NHSE by toggling the topological character of the perpendicular sub-chains, as was first proposed in the cold atom context [316]. When the switch is activated by adjusting the loss parameter, non-reciprocal pumping prevails, driven by non-reciprocity introduced by the switch. Zhang et al. [309] experimentally demonstrated this topological switching using their 3D circuit [Fig. 4e1]. The voltage dynamics across four different unit cells revealed distinct eigenmode distributions, such as skin-topo-skin and skin-topo-topo modes [Fig. 4e2, e3]. The absence or presence of skin modes corresponds to trivial or non-trivial topological modes, which are tuned through relays that switch resistors on or off between specific nodes.
@@ III.2.3 Circuit demonstrations of interacting NHSE phenomena
Despite being classical systems, electrical circuits can also simulate few-body interaction dynamics through mappings between many-body one-dimensional models and single-body multi-dimensional models [317, 48, 318, 319]. Besides their practical value, such mappings can also lend new physical insights [320, 321, 322], such as the interpretation of two-body repulsion as 2-dimensional boundary effects, and 3-body non-Hermitian interactions as feedback mechanisms on a 2-dimensional non-orthogonal non-Hermitian lattice whose GBZ description is notoriously subtle [245]. In particular, effective boundaries can be defined in the Hilbert space even when the lattice has PBCs along each direction. Reverse implementations of such mappings have also enabled small interacting quantum systems to simulate the physics in far larger multi-dimensional lattices [323, 324, 325].
@@ III.2.4 Implementing NHSE without non-reciprocal couplings
Even though paradigmatic NHSE models, i.e., the Hatano-Nelson model [326] almost always contain non-reciprocal (asymmetric) couplings, realizing the NHSE itself does not necessarily require asymmetric couplings. In the so-called reciprocal non-Hermitian skin effect, a reciprocal system is designed such that it contains subsystems harboring equal and opposite NHSE, such that NHSE accumulation occurs in specific subsectors despite the system being reciprocal on the whole. This was first demonstrated by Hofmann et al. [127], who measured eigenmode localization at the edges of their 2D circuit [Fig. 4h1]. A diagonal imaginary hopping within the square plaquette makes the Hamiltonian globally non-Hermitian but still reciprocal; with a basis transform, it can be shown that it contains “hidden” equal and opposite NHSE chains at different transverse momentum sectors or effective flux. This imaginary hopping—realized as a resistor in their model [Fig. 4h2]—results in different hopping probabilities in the presence of this effective flux, resulting in unequal amplitudes in opposite directions [Fig. 4h3].
In two dimensions, a reciprocal 2D lattice can also exhibit nontrivial system size-dependent NHSE in the presence of higher-rank chirality [327], characterized by a non-conserved charge current [328]. First demonstrated in Ref. [327], a higher-rank chiral mode (rank-2) was also demonstrated by Zhu et al. [312] through gain/loss couplings that create an unbalanced chirality in the unidirectional chiral mode flow [Fig. 4i1]. This imbalance causes one specific mode to amplify and hence outlast others, leading to eventual mode localization. The long-lived rank-2 chiral mode dominates the system’s long-time dynamics and is observable in momentum-resolved measurements [Figs. 4i2, i3]. Remarkably, momentum-dependent input excitations lead to mode accumulation at specific driven locations in the lattice [Fig. 4i4]. This enables a versatile implementation of NHSE along a desired direction and position in a fully passive and reciprocal lattice, distinguishing it from the reciprocal skin effect presented in Ref. [127]. (Even in the absence of gain/loss and non-reciprocity in the lattice bulk, non-Hermiticity can emerge through the interaction Hamiltonian in an electromagnetically engineered metamaterial that is inversely designed from a TE lattice [329]. See also Ref. [330], where the NHSE is realized in reciprocal lattices via gauge fields without gain or loss.)
@@ III.2.5 Demonstrating unconventional NHSE phenomena with long-range couplings
The extreme versatility of electrical circuit connections, which can directly couple any two desired nodes independent of their physical distance, makes them ideal for implementing intrinsically long-ranged models [331]. Since the NHSE is already highly non-local in that it leads to dramatic state accumulation in a spatially-distant edge, it is particularly susceptible to the effects of long-ranged hoppings that can change the overall real-space topology of the lattice. For instance, a single non-reciprocal impurity hopping in an otherwise homogeneous Hatano-Nelson chain is known to lead to enigmatic scale-free localization [332, 333, 334].
Recently, Guo et al. [167] introduced the new notion of scale-tailored localization (STL), distinct from conventional NHSE or Anderson localization [335]. The STL is a spectacular demonstration of how singularities in the skin localization length subtly interplays with hopping non-locality, as was implemented via unidirectional couplings with INIC circuit elements, and a long-range coupling controlled by switches [Fig. 4j1]. For this circuit, some of the admittance eigenvalues encircle the unit circle, while others are isolated within it [Fig. 4j2]. The spatial distribution of the corresponding STL eigenstates exhibits varying localization lengths that depend on the long-range coupling distance [Fig. 4j3].
III.3 Circuit applications for non-Hermitian sensing
@@
The primary interest in the NHSE stems from its strong dependence on local perturbations, such as a coupling between the two ends of a chain, which can drastically alter the OBC vs. PBC properties. Furthermore, a non-Hermitian topological lattice exhibits extreme sensitivity to its size due to the exponential growth of NHSE states, as formulated theoretically by Budich and Bergholtz [336], and popularized as non-Hermitian topological sensors [337]. This phenomenon was experimentally demonstrated by Yuan et al. [96] [Fig. 5a1]. The shift in resonance peak frequency occurs through two mechanisms: (i) control of the boundary coupling capacitance through the displacement and rotation of the copper electrodes [Fig. 5a2], and (ii) variations in the number of nodes in the circuit. The impedance resonances exhibit exponential behavior in response to circuit size [Fig. 5a3].
@@
Enhanced response can also be achieved through various mechanisms, such as multiple non-local long-range asymmetric couplings, which lead to an th-order sensitive response, where represents the number of nodes, as introduced in Refs. [95, 97]. Ref. [148] also presents an integrated circuit sensor demonstrating an exponential response with respect to the device size, based on higher-order non-Hermitian topology.
In the previous non-Hermitian sensing experiments, the underlying mechanism is directional amplification, which results in a strongly non-local response. Indeed, exponential sensitivity intuitively requires a non-reciprocal non-Hermitian bulk, achieved through asymmetric couplings [339], which necessitate active elements such as op-amps in INICs [340, 341]. However, Zhang et al. [100] presented a fully passive circuit demonstrating a highly distinct impedance response dependent on the presence of a boundary coupling, even though the circuit uses only resistors as the non-Hermitian components.
The non-local impedance response in this passive circuit is due to the presence of “hidden” coupled Hatano-Nelson chains with opposite amplifications, as seen through a basis rotation [Fig. 5c1]. After the basis transform, the effective momentum space circuit Laplacian at the resonant frequency is expressed as
@@
While this setup may seem to be just another demonstration of the reciprocal NHSE [127] (see Fig. 4h in Section III.2.4), the effective inter-chain coupling is set to be relatively weak compared to the intra-chain terms, such that the NHSE from both chains does not simply cancel in the intended experimental setup. However, due to the exponential sensitivity of the NHSE, they still do cancel for sufficiently long circuits, as a variation of the critical NHSE [299, 236, 300, 301] (refer to the discussion related to Fig. 4b in Section III.2.1). In this model, the circuit exhibits curious size-dependent isolated modes only in the presence of parasitic resistances [Fig. 5c4], when the OBC circuit is sufficiently long. Their spatial profiles are boundary-localized in a unique manner, as shown in the two rightmost plots in Fig. 5c4. Despite being robust against perturbations, such isolated modes do not appear to be protected by any conventional topological invariant, appearing at certain ranges of system sizes.
@@ IV.1 Non-linear phenomena in topolectrical circuits
Introducing non-linear dynamics into a system reveals intriguing phenomena that are unattainable in linear systems, such as higher harmonic generation [140, 342, 343], soliton wave propagation [344, 345, 346, 347], non-linear bands [348], non-linear interface modes [349] as well as various non-Hermitian interplays [350, 351, 352, 353, 354, 208, 355]. Specifically, in the presence of non-linear interactions, coherent collective dynamics can emerge among the autonomous units hosting non-linear oscillators. Of particular interest is the incorporation of a topological mechanism, which facilitates self-organized, protected boundary oscillations [356, 350, 349]. A common approach to realizing non-linear topological phenomena involves arranging non-linear oscillatory elements in a topological configuration, such as by dimerizing an array of oscillators in deference to the well-known SSH configuration.
@@
Lo et al. [359] recently presented a non-linear circuit experiment demonstrating NHSE that emerges depending on the intensity of the driving amplitude. Interestingly, this behavior arises not from asymmetric couplings but from Kerr-type non-linear dynamics [364, 365, 366]. The circuit comprises back-to-back onsite non-linear varactors [Fig. 6h1] assembled into a circuit array inspired by the physics of Bogoliubov modes [367] [Fig. 6h2]. The non-linearity breaks pseudo-Hermiticity, leading to a spectral point gap that in turn gives rise to NHSE localization. The sideband modes [Fig. 6h3], distinct from higher-order harmonics, exhibit localized profiles that depend on the auxiliary drive frequency. Due to the Kerr non-linearity, the NHSE appear only when the input signal amplitude exceeds a certain threshold [Fig. 6h4]. This work highlights how non-linearity can induce the NHSE in classical electrical circuits.
@@
The complex emergent dynamical behavior in non-linear systems can sometimes pose significant challenges for achieving global synchronization [368]. However, Zhang et al. [360] recently demonstrated a Hatano-Nelson circuit hosting on-site non-linear Stuart-Landau oscillators [Fig. 6i1]. Due to the asymmetric couplings in the Hatano-Nelson circuit, Hermiticity is broken, resulting in non-orthogonal eigenmodes. This facilitates more effective communication between modes, ultimately leading to global synchronization. This behavior manifests as collective voltage oscillations observed in the time domain [Fig. 6i2] and in the steady-state voltage profile [Fig. 6i3]. This study highlights how the NHSE can enable global synchronization in non-linear experimental circuits (also refer to Ref. [369]).
@@
A cornerstone of many non-linear systems is their ability to support chaotic dynamics [370, 371, 368, 361]. A paradigmatic chaotic system is the Chua oscillator which, unlike Stuart-Landau oscillators, exhibits irregular and chaotic orbits in some parameter regimes [372]. Mathematically described by the Lorenz system of differential equations, the Chua oscillator possesses a well-known circuit realization known as the Chua circuit [373, 374, 375]. It can be constructed with four passive components and one active component known as the Chua diode [376][Fig. 6j1], which provides the requisite non-linear negative resistance. This non-linearity is evident in its non-linear current-voltage () characteristic [Fig. 6j2], and typical chaotic behavior includes the famed double-scroll chaotic attractor [Fig. 6j3].
@@
Of particular interest is how such chaotic behavior interplays with lattice band structure dynamics, such as when multiple non-linear chaotic oscillators are coupled to form an array. In Ref. [361], Sahin et al. investigated the consequences of coupling Chua oscillators in a topological SSH-like lattice structure [Fig. 6j4]. Since band topology is only rigorously defined in linear systems, one might intuitively expect the non-linearity, and especially its chaotic dynamics, to erode the topological localization. However, it turns out that topological edge localization persists robustly well into the non-linear regime, with the amplitudes of the chaotic scroll being very different at the bulk and edge oscillators [Fig. 6j5] only in the black parameter space region [Fig. 6j6]—the region extrapolated from the topological non-trivial parameter region defined in the linear limit. This extrapolation can be performed by tuning the - characteristic parameter of the Chua diode [Fig. 6j7], and explicitly showcases how the topological edge oscillation suppression robustly carries over from the linear into the non-linear regime.
@@
By externally perturbing the voltage oscillations by injecting current at the left edge node, one further observes that the chaotic oscillations, which are typically highly sensitive to external perturbations, persist robustly only when in the topologically non-trivial regime. In the regime connected to the trivial phase, the injected current reduces the chaotic double-scroll portraits to non-chaotic limit-cycle oscillations [Fig. 6j8]. However, in the non-trivial phase, the same current injection only disturbs the edge node oscillations, leaving the bulk oscillations mostly intact [Fig. 6j9]. This robustness is captured as a phenomenon termed “topologically protected chaos”, where chaotic dynamics counterintuitively become more robust by the remnants of topological localization deep inside the non-linear regime [361].